OpenStudy (anonymous):

Mindy, Kristie and Amy are trying to solve this equation: 2x + 4(x + 1) - 3 = 15. Mindy says the first step is to combine 2x and -3. Kristie says the first step is to add 3 to both sides. Amy says the first step is to distribute and multiply both terms in the parentheses by 4. Who is correct? A. Mindy B. Kristie C. Amy D. no one

OpenStudy (anonymous):

hold on i wanna try an figure it out

OpenStudy (anonymous):

C. like the saying PEMDSA

hartnn (hartnn):

YES

OpenStudy (anonymous):

C. is sorrect following the order of operation PEMDAS as you stated

OpenStudy (anonymous):

PARENTHESIS EXPONENT MULTIPLY DIVIDE ADITION SUBTRACT :)

OpenStudy (anonymous):

THanks!! :)

hartnn (hartnn):

welcome ^_^

OpenStudy (anonymous):

The girls are working on another problem together. They are trying to solve the equation: 2(x + 5) - 4 = 36 Mindy says the answer is x = 15. Kristie says the answer is x = 17.5 Amy says the answer is x = 13 Who is correct? A. Mindy B. Kristie C. Amy D. no one

hartnn (hartnn):

2(x + 5)=?

OpenStudy (anonymous):

2x+ 10 =12x?

hartnn (hartnn):

no, its just 2x+10 can't write 10 as 10 x so now what is 2x+10-4

hartnn (hartnn):

combine like terms, 10 and -4 are like, 2x is different

OpenStudy (anonymous):

6 an d 2x

hartnn (hartnn):

so u get 2x+6 =36 know what next step will be ?

OpenStudy (anonymous):

getting rid of 2 from x?

hartnn (hartnn):

getting rid of 6 by subtracting 6 from both sides

OpenStudy (anonymous):

-4?

OpenStudy (anonymous):

or 4

hartnn (hartnn):

2x+6-6 = 36-6 what is 6-6 =? and 36-6= ?

OpenStudy (anonymous):

0 and 30

hartnn (hartnn):

so u have 2x=30 now get rid of 2 by ?

OpenStudy (anonymous):

taking away 2 from both sides/

hartnn (hartnn):

dividing 2 , if that what u meant x=30/2 = ?

OpenStudy (anonymous):

my bad dividing?

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

30/2=15 2/2=0

OpenStudy (anonymous):

A.?

hartnn (hartnn):

2/2 = 1 so u have 1x=15 that is x=15

hartnn (hartnn):

yes, A

hartnn (hartnn):

understood the steps ?

OpenStudy (anonymous):

yes long but i do now thank you :)

hartnn (hartnn):

welcome ^_^